OpenStudy (anonymous):

factor 1/4x^2-1/4 comepletely

OpenStudy (anonymous):

\[1/4x^{2}-1/4\]

OpenStudy (anonymous):

1/4x2−1/4 = 1/4(x2−1) =1/4 (x-1) (x+1)